# PEARL: Peer-Enhanced Adaptive Radio via On-Device LLM

### v1

@@ 1 Introduction
In D2D communication, link performance depends not only on the local device state but also on peer conditions such as battery level or device type. For example, when the publisher has ample energy but the subscriber is low on battery, an energy-preserving configuration may be preferable even at the cost of slightly higher latency. Designing a learning system that accounts for such asymmetric, two-sided context remains an open challenge. Moreover, efficient on-device deployment requires parameter-efficient fine-tuning (PEFT) strategies (Bucher and Martini, 2024; Han et al., 2024), but it is unclear how different PEFT modules compare in this setting, or how to ensure that training signals faithfully reflect heterogeneous application and device constraints.
@@ 2 Problem Formulation
Objective. The objective is to minimize link latency and device energy consumption while meeting application-specific requirements. Both local (publisher) and peer (subscriber) context guide parameter selection. For example, when the subscriber has low battery, the model should prefer energy-preserving modes even at the cost of slightly higher latency.
@@ 3.1 System Architecture
PEARL builds on a pre-trained large language model (LLM) as a general-purpose feature encoder, and adapts it into a lightweight classifier for cross-layer parameter selection (Wu et al., 2024). Context features from both publisher and subscriber devices (e.g., battery levels, device type, time, application) are embedded into a structured prompt and passed to the frozen LLM backbone. A task-specific classification head is attached on top, predicting one of the discrete parameter tuples $(\texttt{PerformanceMode},\texttt{AccessCategory})$.
@@ 4.1 Setup
Baselines. We evaluate our proposed framework for context-driven cross-layer optimization under the PEARL and PEARL-Lite variants introduced in Section 3.1. For comparison, we also include: (i) an oracle, which selects the global maximizer using ground-truth latency and energy; (ii) LoRA-only (LoRA); (iii) a rule-based baseline (Rule), which selects WA parameters heuristically based only on application type (see Table 8 in Appendix E); and (iv) two fixed baselines, Fixed (RT/IV) (fix-RT/IV, $(\texttt{realtime},\texttt{interactiveVoice})$) and Fixed (Bulk/BG) (fix-Bulk/BG, $(\texttt{bulk},\texttt{background})$).
@@ 4.3 Impact of Peer Context Information
Table 1: Impact of including peer information in the context. In the aggregated scenario, models trained w/ peer info achieve consistently higher objective scores than those w/o peer info. In the cooperative case (subscriber low battery), adding peer info reduces energy by up to $\sim$16% while maintaining latency, highlighting where peer awareness is most impactful.
@@ Appendix D Data Collection
The simulation dataset is constructed from real WA logs, combined with synthetic application usages defined in Fig. 8. To collect logs, we developed an iPadOS app that establishes a WA connection between two devices, simulates their interactions, and records context information. The app was deployed on two iPad Pro devices (Table 4), where we executed 16 sessions of data collection, each consisting of 2,000 consecutive logs. Each session corresponds to one specific scenario defined by time of day and device battery levels.
@@ Appendix E Baseline Details (Rule)
This design makes Rule adaptive only to recent application history, without considering device battery states or other contextual factors. It thus serves as a lightweight non-learning baseline for comparison against context-aware methods such as PEARL and PEARL-Lite.

### v2

@@ 1 Introduction
In D2D communication, link performance depends not only on the local device state but also on peer conditions such as battery level. For example, when the publisher has ample energy but the subscriber is low on battery, an energy-preserving configuration may be preferable even at the cost of slightly higher latency. Designing a learning system that accounts for such asymmetric, two-sided context remains an open challenge. Moreover, efficient on-device deployment requires parameter-efficient fine-tuning (PEFT) strategies (Bucher and Martini, 2024; Han et al., 2024), but it is unclear how different PEFT modules compare in this setting, or how to ensure that training signals faithfully reflect heterogeneous application and device constraints.
@@ 2 Problem Formulation
Objective. The objective is to balance link latency and device energy consumption while satisfying application-specific requirements. Both local (publisher) and peer (subscriber) context guide parameter selection. For example, when the subscriber has low battery, the model should prefer energy-preserving modes even at the cost of slightly higher latency.
@@ 3.1 System Architecture
PEARL builds on a pre-trained large language model (LLM) as a general-purpose feature encoder, and adapts it into a lightweight classifier for cross-layer parameter selection (Wu et al., 2024). Context features from both publisher and subscriber devices (e.g., battery levels, time, application) are embedded into a structured prompt and passed to the frozen LLM backbone. A task-specific classification head is attached on top, predicting one of the discrete parameter tuples $(\texttt{PerformanceMode},\texttt{AccessCategory})$.
@@ 4.1 Setup
Baselines. We evaluate our proposed framework for context-driven cross-layer optimization under the PEARL and PEARL-Lite variants introduced in Section 3.1. For comparison, we also include: (i) an oracle, which selects the configuration that maximizes the context-aware reward derived from ground-truth latency and energy — effectively representing the optimal balance between latency and energy consumption under the defined application constraints; (ii) LoRA-only (LoRA); (iii) a rule-based baseline (Rule), which selects WA parameters heuristically based only on application type (see Table 8 in Appendix E); and (iv) two fixed baselines, Fixed (RT/IV) (fix-RT/IV, $(\texttt{realtime},\texttt{interactiveVoice})$) and Fixed (Bulk/BG) (fix-Bulk/BG, $(\texttt{bulk},\texttt{background})$).
@@ 4.3 Impact of Peer Context Information
Table 1: Impact of including peer information in the context. In the aggregated scenario, models trained w/ peer info achieve consistently higher objective scores than those w/o peer info. In the cooperative scenario—where the subscriber is low on battery and the publisher adapts its configuration to assist—adding peer info reduces energy consumption by up to $\sim$16% while maintaining reasonable latency, highlighting where peer awareness provides the greatest benefit.
@@ Appendix D Data Collection
The simulation dataset is constructed from real WA logs, combined with synthetic application usages defined in Fig. 8. To collect logs, we developed an iPadOS app that establishes a WA connection between two devices, simulates their interactions, and records context information. The app was deployed on two iPad Pro devices (Table 4), where we executed 16 sessions of data collection, each consisting of 2,000 consecutive logs. Each session corresponds to one specific scenario defined by time of day and device battery levels. Battery levels were sampled uniformly across high, medium, and low states to ensure balanced representation of energy conditions.
@@ Appendix E Baseline Details (Rule)
This design makes Rule adaptive only to recent application history, without considering device battery states or other contextual factors. It thus serves as a lightweight non-learning baseline for comparison against context-aware methods such as PEARL and PEARL-Lite. Future extensions could incorporate a simple “battery-safe” heuristic—e.g., switching to energy-preserving modes below 20% battery—to better reflect real-world smartphone behavior.